OpenStudy (anonymous):

Given the two figures are similar, calculate the value of w. file:///Users/jellybean/Desktop/Geometry_6_Exam_Final_NK_051209_q80.gif

geerky42 (geerky42):

We cannot access your file on your computer. Post it on internet.

geerky42 (geerky42):

\[\dfrac{2}{2.5} = \dfrac{3}{w}\]Do cross multiply. \[2w = 7.5\]\[w = \dfrac{7.5}{2} = \boxed{3.75}\] Is this clear? Can you find the value of v?

OpenStudy (anonymous):

Yes that's clear! So would v be 4.24?

geerky42 (geerky42):

Close, it's 4.16. \[\dfrac{2}{2.5} = \dfrac{v}{5.2} \Rightarrow v = \dfrac{2\times 5.2}{2.5}\]

OpenStudy (anonymous):

Ohhh okay thank you so much!